Mangum wilson eelco hills hunter wootten richer nyman emerson nikolic Marconi Lucas

four ntennas at high site Monthly reports Astroholography high res maps done on Saturn. These reveal we are measuring most of surface error resolution about .5m in 96x96 maps. Features have some elements in common. grav residuals and thermal effects expected. Some common features similar to DV02 after return to osf. These are 17-18 microns, within spec but predictions had posited lower errors. We need to understand this. DV02 has just rearrived but still shows some errors. Some of bottom and top structure due to illumination offset with this receiver. We have known of this and attributed it to misalignment of receiver optics. PM03 shows an offset but different, resolution good enough to see opt hole. stigmatism has been removed from these maps. One still sees a fourfold pattern however. This was known in melco antennas as a temp-functional effect. pm03 has not had all improvements to thermal control made. JR: illumination? HR we use external optics in B3 and the procedure for alignment needs work, not even done on some early FEs. That needs improvement. TH: how long rh two hours, a lot of that software overhead. rh: mode testing seems to be going very well. These were basic tests of es modes. phase stability. The figure here is based on a report from Morita who is not available today. Not sure if periodic phase glitches affected this data. Cloud monitoring. bn: use WVR to measure opacity not associated with water vapor; this correlates very well with visual estimates. Variable cloud cover affects phase correction and is a useful ancillary information to interpret WVR. rh: Software v 7.1 went pretty smoothly. This provides automatic control of focus. Some thingw didn't make the release. subscan scheduling didn't make it for example. Probably in next round a month away. With five antennas we saw the correlations not coming out right. At five the signals come through a separate station bin and problems showed up but we believe those are now ok. Pointing is still a problem, some element is due to inclinometer drifts on hour scales on dv01 particularly. This will be changed at July. These appear to be thermal effects. Availability of antennas is still a problem; systems do not always return to a normal state. We have two antennas now, maybe another one tonight. Antenna people are having to put a lot of effort into recovering antennas also. New software allows 180 deg switching which allows us to do B9, some testing of frequency offsetting done last night. RL: This was tested last nite and worked perfectly.

Gianni Marconi has joined the team. NAASC: Leroy and Schnee hired, third position will be readvertised. Nyman: interviewed, will make 4 job offers for end of year. DM interviews upcoming Lead person for proposal handling team out. rh: there are posts available. Please let people know.

PS Report on AAS meeting, workshop CASCA 200 people, a session on preparing for alma with a demo of OT, both had ~3 dozen people. JR: went to B5 meeting, many good science talks. Nothing surprising but a key role for molecular and C+ lines. The SOC will write a short science case.

ASAC--still await a response from the project to appear in the next few days.

CRE on B6 has not been signed off. The Board does have to make a determination. That discussion would happen soon. A revision dated 21 May is on edm. this sets high limits across the band; this is the best FE can guarantee. The do not wish to change the feed horns any time soon. If main problem is due to feed horns it should be stable, so we may be able to calibrate it. This will affect tp accuracy unless we always make a full stokes analysis. I think we don't have any real choice but to do this. We need to press on the project to improve the design. rh will try to organize a discussion. Spurious responses. DTE: 1st LO mixes down to 4-12 ghz, second then mixes this down to the freq where it is digitized. At the top end of 4-12 this could be normally the lsb from signals at top end of 4-12 ghz if. Owing to filter imperfections signals from outside the band will appear in the usb of the if. For a small range of lo2 frequencies seen in the document which will show that we could get in the extreme case a signal 8db down. Appearance in data--a freq outside the if unexpected. the phase of the unwanted signal will have its phase reversed so that in maps there will be a mirror image in them. So unexpected signal in data, and inappropriate place in sky will show up. This can be complex to diagnose. But this is from a very specific part of the band, which may be rare in occurrence. Remedies are a software fix, to keep LO2 from the bad frequencies. Hardware fixes also possible. What does science group prefer? Science could ask for an assessment A second problem. LO1 comes from multiplied yig signal. 12 GHz is top end of the if range, some signals above that can leak and be aliased back through so that with certain LO1 freqs a strong CW signal can get through. This could be mitigated with LO offsetting etc. The same hardware cure proposed above could cure this also. How worried is the science group about this. We don't know what will be entailed. This will affect total power and if. RL: will the first problem not be removed by LO offsetting? DTE: Yes, that is discussed in Peter's document. phase switching would not reject it. Still affects total power. rh: in principle if phase switching in 2nd lo taking it out in correlator would be a problem. b9 we need it in the 1st LO for sideband separation. I don't see this as very urgent. Do we need more experiments? Is the analysis complete? I have a background worry that this is a large effect occurring in exceptional circumstances. A lower level effect may be present under a wider range of circumstances. In the design these things should have been thought about but perhaps not all paths for spurious signals were contemplated. It may be we could solve this problem by the software fix. DTE: There have been more experiments since Peter's document was written we do need. RH: this feature seemed to have a significant width, not clearly a CW line. DTE: May still be a problem.

OT--subset to be distributed. TH--this will be a solution for the EVLA also. Recommend the shorter list of lines. Nyman--ph 1 manager. testing in final stages. getting reports. ot was used to submit proposals for some time to test the server etc. We did have some problems. We received about 200 proposals. There were lots of details which needed to be sorted out. A load test of 200 proposals submitted in 15 minutes seemed to work pretty well. Ph1 manager things seemed to go pretty well. Tech and sci assessment, also simulation of the prc meetings. Then ph 2 submission, pt being used to track status. Next test in Aug and Sept with updated tools in prep for readiness review in October. One can select four identical bands in OT those this seems of little use. rh: overall this sounds like it works pretty well.

Calibration. There was a broken link for the met station.

location of met station. proposed location near a91 on lhs side of map attached to agenda. Eduardo Donoso noted that if there we would have to wait for a year owing to the power and fiber implementation strategy, and second suggestion near A88 has a possibility. We can have a location there soon with power and commo, in a month or so. We could move later. The A88 does not fall west of the central array, is fairly near the road which is a major traffic road. Waiting 1.5 years will be a non starter. thermal sounder would go here also. also wvr, also the 11 ghz phase monitor.

We may wish to move but let us proceed with this. We now. WVR does need to be inside an enclosure. AI: Al to put this in. Power and fiber on slab.

-- AlWootten - 2010-06-14
Topic revision: r3 - 2010-08-16, AlWootten
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback